Units Materials Conversion Work in process inventory, beginning 10,000 $5,700 $20,400 Units started in process 90,000...

Units

Materials

Conversion

Work in process inventory, beginning

10,000

$5,700

$20,400

Units started in process

90,000

Units transferred out

84,000

Work in process inventory, ending

8,000

Costs added during the month

$52,800

$76,400

The beginning work in process inventory was 80% complete with respect to materials and 60% complete for Conversion. The ending work in process inventory was 75% complete with respect to materials and 50% complete with respect to conversion costs.

5) What are the equivalent units for Materials and Conversion?

a). Materials – 84,000, Conversion – 84,000 c.) Materials – 98,000, Conversion – 96,000 b.)Materials – 92,000, Conversion – 90,000 d.) Materials – 90,000, Conversion – 88,000

6) What is your cost per Equivalent Unit based on what you calculated in question 5 above?
Materials ____________ + Conversion ______________ = Total Cost ___________________

7) What is your recalculated cost of Ending Inventory?

  1. a. $8,300 b. $155,300 c.$14,000 d.154,000 .

  2. 8) What is your cost of Goods Transferred?

  3. a. $157,500 b.147,000 c. $14,000 d. $155,300

Prepare a Cost Reconciliation report for the information related to problems 5 through 8.

Please show work for each problem!

Those units are all i was given
Answer #1
Statement of Equivalent production
Particulars Direct Materials % of completion Equivalent material Conversion % of completion Equivalent Conv Total
Units transferred out 84,000 100% 84,000              74,000 100% 84,000
ending WIP                8,000 75%                    6,000                8,000 50%                    4,000
Equivalent Units                 90,000                 88,000
Opening WIP Costs                    5,700                 20,400      26,100
Current Month Costs                 52,800                 76,400    129,200
Total costs                 58,500                 96,800    155,300
Cost per Equivalent units                            0.65                            1.10                 1.75
Cost of ending inventory (6,000 x 0.65) + (4,000 x 1.10) = 3,900 + 4,400 = 8,300
Cost of goods transferred = 84,000 x 1.75 = 147,000   

Cost reconciliation

Cost to be accounted for
Beginning work in process 26,100
Cost added 129,200
Total Cost to be accounted for 155,300
Cost accounted as
Cost of unit transferred out 147,000
Cost of ending work in process 8,300
Total cost accounted as 155,300